Grade 11Mechanics

Ball A is thrown up vertically with speed 10m/s and at the same instant another ball B is released from rest at hieght h. At time t ,the speed of A relative to B is
(a) 10m/s
(b) 10-2gt
(c) √10^2-2gh
(d)10-gt

Khushboo
8 Years agoGrade 11

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
Dear student
 
 
V of a= 10+gt. ( Taking upward direction as +ve)
V of b= 0+(-g)t = -gt
We know that relative v = v of a- v of b
But here both the velocities are opposite to each other in directions so take one of them as -ve ( here taking v of b as -ve)
So relative velocity= v of a -(-vof b)
= 10+gt +(-gt)
=10m/s
 
Hope it helps
 
 
Regards
Arun (askIITians forum expert)